when a ball fall from rest it crashes a glass plate after travelling  6 sec and its velocity is decreased by 2/3 then after 2 sec it reaches the ground then find the distance travelled by ball?

Grade:

2 Answers

Chetan Mandayam Nayakar
12 years ago

distance travelled=(g/2)*62+((1/3)g*6)*2+(g/2)*22

Nikhil panduranga
12 years ago

I dont know whether this answer is right or wrong....but,i've still attempted this....

IN THE FIRST LEG OF THE JOURNEY,,ie FROM THE INITIAL POSITION TO THE GLASS PANE .THE DATA IS u=0;v=x(say)ms^-1;t=6secs;

v^2=u^2+2aS,since u=0;   v^2/2a=S 

a=v/t=.....................

IS THE ANSWER 11x/3 m.......

If thats right i will tell u how it solved this....